Circular No. 8809

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1280 SCORPII R. J. Rudy, D. K. Lynch, and R. W. Russell, The Aerospace Cor-

poration; C. E. Woodward, University of Minnesota; and J. Liebert and M. Cushing, University of Arizona, report 0.8- to 5-micron spectroscopy of V1280 Sco from Feb. 14 and 16 UT, during the nova’s rise to maximum (Feb. 17.75 UT), using the SPEX instrument on the 3-m reflector of the Infrared Telescope Facility: “The spectrum was dominated by the absorption lines of hydrogen, and there was a pronounced jump due to the Paschen discontinuity. Emission lines were almost entirely absent, with Paβ and Brα just beginning to exhibit hints of emission and P-Cyg behavior. There were also numerous absorptions due to low-lying but non-resonant transitions of neutral nitrogen and carbon. The infrared continuum, with the exception of the Paschen and Brackett discontinuities, was well matched by a 5300-K blackbody. Between the two dates, the H I absorption lines’ equivalent widths decreased in strength by about a factor of two.”

Visual magnitude estimates: Feb. 16.735 UT, 4.3 (A. Pearce, Nedlands, W. Australia); 17.316, 4.2 (W. Souza, Santos, Brazil); 18.205, 4.7 (A. Kammerer, Malsch, Germany); 19.145, 4.6 (J. de S. Aguiar, Campinas, Brazil); 19.416, 4.1 (R. Donner, Marblehead, MA, USA); 20.140, 4.7 (Aguiar).

COMET C/2007 D2 (SPACEWATCH) An object discovered by Spacewatch (discovery observation tabulated

below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to show cometary appearance on CCD images taken by M. Tichý and J. Tichá (Kleť 1.06-m reflector; Feb. 18.95 UT; object diffuse with coma diameter 12′′ and faint tail in p.a. 90^{o}) and by L. Buzzi (Varese, Italy, 0.60-m f/4.64 reflector; Feb. 20.1; object clearly non-stellar; stacked images totalling 21 min of exposure time show a compact coma almost 9′′ wide with an extension 12′′ long in p.a. ∼ 100^{o}).

2007 UT  α_{2000}  δ_{2000}  Mag.
Feb. 17.45924  12^{h}24^{m}27.^{s}87  −0^{o}21′57′′.1  20.1

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-D27.

T = 2006 Nov. 25.670 TT  ω = 66.^{o}245
Ω = 296.918 }2000.0
q = 1.26149 AU  i = 178.620
2007 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green